# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v24

@@ 3.24 Sinhala Braille
The work by de Silva and Liyanage [452] uses KNN, SVM, and a simple ANN system to recognize Sinhala Braille text. Vithanage [652] also claims to have created a conversion engine to easily convert the Braille text into the corresponding Sinhala text. The study by Madubashana [653], the study by Ariyarathna et al. [654] as well as the study by Weerasinghe [655] also focused on creating an automated Braille to Sinhala recognition system.  Anuradha and Thelijjagoda [363] uses OCR on Sinhala and English Braille documents on which they then run a machine translation system in order to convert them into voice.
@@ 3.26 Sinhala-English Code-Mixing
As discussed in Section 3.10Rathnayake et al. [253] used adapter-based [254, 255, 256, 257, 258, 259] fine-tuning on XLM-R [260], for classifying code-mixed and code-switched Sinhala text. Hettigoda [263] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [283] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [265] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [611] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [451] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [668] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by Udawatta et al. [669] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection. Uthpala and Thirukumaran [285] claims to have created a sentiment-annotated Sinhala-English code mixed dataset using comments from YouTube videos. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
@@ 3.27 Language Acquisition and Evaluation
Sandathara et al. [670] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [671] assists children to learn to speak without stuttering. Nethmi et al. [672] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [673] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [674] uses SVM and LSTm to identify phonological disorders from Sinhala speech. Francis et al. [675] claims to use Sinhala OCR and PoS tagging in their effort to build the SinhaLearn, a system for improving Sinhala Proficiency of grade 5 scholarship students. Praveen et al. [362] also claims to have created a web-based education system that helps grade 5 scholarship students who are learning the Sinhala language. They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system.
@@ 3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [676]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [107] has published two Sinhala BERT models: SinBERT-large150150150https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small151151151https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo152152152https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [677] on the OSCAR [155, 156] dataset. Demotte et al. [272] have released the word embeddings153153153https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code154154154https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer. Gurgurov et al. [678] included Sinhala as one of their 87 low-resource languages for which they provided word embeddings in their LowREm repository. They credit Demotte et al. [272] as their source for Sinhala embeddings.
@@
Wickramasinghe and de Silva [148] used Sinhala-English dictionary data [147] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub155155155https://bit.ly/3t3SKu7. Subsequent work by Rathnayake et al. [679] also attempts BLI for Sinhala English but uses UVecMap (Unsupervised VecMap [680]) as the baseline. They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [681] archives the best result of 29.49 for En-Si FastText embedding.
@@
Üstün et al. [171] introduced Aya, a massively multilingual generative language model156156156https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159]. Jayakody and Dias [682] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude157157157https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o158158158https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)159159159https://llama.meta.com/docs/get-started/ and Mistral 7B160160160https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2161161161https://falconllm.tii.ae/ and Gemini162162162https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala. Chang et al. [683] introduced Goldfish, a collection of monolingual language models163163163https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109]. The work by Ayesha et al. [684] used a fined-tuned Llama 3.1 (8B) model [685] to predict student performance by analyzing textual entries created by students in Sinhala and English.
@@
Vayani et al. [164] analyzed the performance of a number of LLMs over 100 languages that included Sinhala. The basic dataset was created by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165]. This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category. This dataset, which they named ALM-Bench164164164https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available. They made a number of observations in this aspect. First, they observed that for the Q&A task, the performance of GPT-4o [166] for Sinhala is better than Qwen2-VL [686]. Next, they report that including images in the question further improved Sinhala results of GPT-4o by 38.7%. In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding.
@@ 3.29 Text Simplification
The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [687] included the NSINA [108] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data165165165https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data166166166https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [688] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [689] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [690] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [691] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
@@ 3.30 Tokenizing
The study by Petrov et al. [692] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [693] and mT5 [336, 693] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [694]. The work by Velayuthan and Sarveswaran [695] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala. Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available167167167https://github.com/Buddhilive/sltk. However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words. While the authors have made the dictionary publicly available168168168https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries. So while it may be adequate to handle Sri Lanka-centric news title data set169169169https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
@@ 3.31 Detection or Management of Mental Illnesses
Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [668] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [696] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [697] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. Fazal and Farook [698] compared multiple algorithms to detect depression in Sinhala English code mixed social media posts and reported that ExtraTreesClassifier [699] obtains the best results. Welarathna et al. [413] used CNNs to classify emotions of Sinhala speech by Autistic children. They handle the following emotions: sad, disgust, surprise, neutral, happy, calm, fear, and angry. Jayawardena et al. [700] used a NAO Robot [701] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children.

### v25

@@ 3.24 Sinhala Braille
The work by de Silva and Liyanage [467] uses KNN, SVM, and a simple ANN system to recognize Sinhala Braille text. Vithanage [680] also claims to have created a conversion engine to easily convert the Braille text into the corresponding Sinhala text. The study by Madubashana [681], the study by Ariyarathna et al. [682] as well as the study by Weerasinghe [683] also focused on creating an automated Braille to Sinhala recognition system. Dampalessa and Ragel [684] used a data set of 230 characters to train a model to transliterate Braille letters into Sinhala text.  Anuradha and Thelijjagoda [376] uses OCR on Sinhala and English Braille documents on which they then run a machine translation system in order to convert them into voice. Udana et al. [685] introduces a prototype tactile Braille display for Sinhala. They claim to be able to convert digital Sinhala text into Braille using an electromechanical process that also utilises a CNN.
@@ 3.26 Sinhala-English Code-Mixing
As discussed in Section 3.10Rathnayake et al. [255] used adapter-based [256, 257, 258, 259, 260, 261] fine-tuning on XLM-R [262], for classifying code-mixed and code-switched Sinhala text. Hettigoda [265] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [287] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [267] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [632] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [466] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [698] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by Udawatta et al. [699] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection. Uthpala and Thirukumaran [289] claims to have created a sentiment-annotated Sinhala-English code mixed dataset using comments from YouTube videos. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. Senanayaka et al. [700] proposes SingRAG, a RAG and translation framework built on LLaMA that converts code-mixed Sinhala to English for processing and then converts whatever the English results is of the attached backend to code-mixed Sinhala before sending it back to the user.
@@ 3.27 Language Acquisition and Evaluation
Sandathara et al. [701] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [702] assists children to learn to speak without stuttering. Nethmi et al. [703] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [704] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [705] uses SVM and LSTm to identify phonological disorders from Sinhala speech. Francis et al. [706] claims to use Sinhala OCR and PoS tagging in their effort to build the SinhaLearn, a system for improving Sinhala Proficiency of grade 5 scholarship students. Praveen et al. [375] also claims to have created a web-based education system that helps grade 5 scholarship students who are learning the Sinhala language. They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system. Weerasinghe et al. [707] introduced LexiLearn, a gamified system intended to help Sinhala learners with dyslexia. The survey by Perera and Sumanathilaka [708] discusses the lack of annotated datasets and NLP resources to cater for Sinhala-speaking adults with dyslexia.
@@ 3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [709]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [107] has published two Sinhala BERT models: SinBERT-large155155155https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small156156156https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo157157157https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [710] on the OSCAR [155, 156] dataset. Demotte et al. [276] have released the word embeddings158158158https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code159159159https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer. Gurgurov et al. [711] included Sinhala as one of their 87 low-resource languages for which they provided word embeddings in their LowREm repository. They credit Demotte et al. [276] as their source for Sinhala embeddings. Gurgurov et al. [712]160160160https://github.com/d-gurgurov/Knowledge-Driven-Adaptation-LLMs argued that mLMs such as mBERT and XLM-R are better suited for low-resourced languages such as Sinhala as opposed to LLMs and suggested parameter-efficient adapter-based methods for adapting such mLMs.
@@
Wickramasinghe and de Silva [148] used Sinhala-English dictionary data [147] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub161161161https://bit.ly/3t3SKu7. Subsequent work by Rathnayake et al. [713] also attempts BLI for Sinhala English but uses UVecMap (Unsupervised VecMap [714]) as the baseline. They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [715] archives the best result of 29.49 for En-Si FastText embedding.
@@
Üstün et al. [171] introduced Aya, a massively multilingual generative language model162162162https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159]. Jayakody and Dias [716] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude163163163https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o164164164https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)165165165https://llama.meta.com/docs/get-started/ and Mistral 7B166166166https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2167167167https://falconllm.tii.ae/ and Gemini168168168https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala. Hansaka and Sumanathilaka [717] also claims to have conducted a review on fine-tuning approaches for LLMs such as GPT-3, GPT-4, and LlaMA in the context of Sinhala. Chang et al. [718] introduced Goldfish, a collection of monolingual language models169169169https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109]. The work by Ayesha et al. [719] used a fined-tuned Llama 3.1 (8B) model [720] to predict student performance by analyzing textual entries created by students in Sinhala and English. Tang et al. [721] evaluated the susceptibility of LLMs being tricked into answering illegal questions in low-resource languages when the same questions would have been successfully rejected had they been asked in a high-resourced language. They provide the framework for testing with the results on eight low-resourced languages, which include Sinhala.
@@
Vayani et al. [164] analyzed the performance of a number of LLMs over 100 languages that included Sinhala. The basic dataset was created by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165]. This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category. This dataset, which they named ALM-Bench170170170https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available. They made a number of observations in this aspect. First, they observed that for the Q&A task, the performance of GPT-4o [166] for Sinhala is better than Qwen2-VL [722]. Next, they report that including images in the question further improved Sinhala results of GPT-4o by 38.7%. In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding.
@@ 3.29 Text Simplification
Ranathunga et al. [723] introduced SiTSE, a Sinhala text simplification dataset and the associated benchmark. They claim that the data set is publicly available on github171171171https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation; however, currently it resolves to a 404. The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [724] included the NSINA [108] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data172172172https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data173173173https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [725] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [726] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [727] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [728] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
@@ 3.30 Tokenizing
The study by Petrov et al. [729] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [730] and mT5 [349, 730] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [731]. The work by Velayuthan and Sarveswaran [732] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala. Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available174174174https://github.com/Buddhilive/sltk. However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words. While the authors have made the dictionary publicly available175175175https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries. So while it may be adequate to handle Sri Lanka-centric news title data set176176176https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
@@ 3.31 Detection or Management of Mental Illnesses
Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [698] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [733] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [734] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. Fazal and Farook [735] compared multiple algorithms to detect depression in Sinhala English code mixed social media posts and reported that ExtraTreesClassifier [736] obtains the best results. Welarathna et al. [428] used CNNs to classify emotions of Sinhala speech by Autistic children. They handle the following emotions: sad, disgust, surprise, neutral, happy, calm, fear, and angry. Jayawardena et al. [737] used a NAO Robot [738] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children. The system by Kulasooriya et al. [739] used SVM and simple ANN to detect psychological stress in Sinhala speech.
@@ 3.32 Sinhala Text Generation
3.33 Question Answering Systems (Q&A)
Dissanayake and Hettige [749] implemented a question-and-answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs. Jayakody et al. [750] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha.
Romero et al. [161] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala. They have publicly released both their data177177177https://huggingface.co/datasets/afaji/cvqa and leader board178178178https://eval.ai/web/challenges/challenge-page/2305/overview. Vayani et al. [164] created a dataset ALM-Bench179179179https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web. They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [166] is superior to other tested LLMs. Wansekara and Jayasekara [751] developed a visual Sinhala question and answer system using YOLO for object identification and a BERT model trained on a Sinhala corpus for Sinhala text. Dhanawardhana et al. [752] frame their work just as a multi-model deep learning approach. But in actuality, their lost-and-found application is utilising techniques that are relevant to Sinhala VQA. The system created by Dasanayaka et al. [634] is claimed to be capable of first, generating radiology reports for orthopantomography (related to dentistry) using Llama on captions generated by a Blip-2-based caption generator on dental panoramic tomography (DPT) images and then serve a Sinhala chatbot where the patients may ask questions on the said report and get answers.
3.34 Language Identification (LID) Task
Burchell et al. [146] show that they obtain better results for the LID task for most of the languages compared to NLLB [134]. However, for Sinhala the results by Burchell et al. [146] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Kargaran et al. [753] introduces a tool named GlotScript which is used for written LangID it achieves above 80% accuracy for Sinhala.